OpenStudy (anonymous):

calculate the circumference of a CIRCLE area is equal to the sum of of the area of the circles with the diameter 24cm,32cm and 96cm?

OpenStudy (anonymous):

@Jhannybean

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

@mathmath333 @M,aliha @Mikayla_Rose

OpenStudy (anonymous):

@hartnn

OpenStudy (jhannybean):

Hmm... for starters.\[A=\pi r^2\]\[C= 2 \pi r\]

OpenStudy (jhannybean):

\[A_T = \pi ((r_1)^2+(r_2)^2 +(r_3)^2)\]

OpenStudy (jhannybean):

\[r=\frac{D}{2}\]

OpenStudy (jhannybean):

So find all your radiuses first.

OpenStudy (anonymous):

144 , 256, 2304?

OpenStudy (anonymous):

so 144 + 256 + 2304 so 2704

OpenStudy (anonymous):

@Jhannybean

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@M,aliha

OpenStudy (anonymous):

@ganeshie8

OpenStudy (jhannybean):

Alright, so \(A_T = 2704\pi\)

OpenStudy (anonymous):

now?

OpenStudy (jhannybean):

now you resolve for \(r\)\[r=\sqrt{\frac{A_T}{\pi}}\]Plug this \(r\) into the formula for circumference and find it.

OpenStudy (anonymous):

52/7?

OpenStudy (anonymous):

i meant sqrt 52/7

OpenStudy (jhannybean):

\[C=2\pi\sqrt{\frac{A_T}{\pi}}\]

OpenStudy (anonymous):

ummm pi gets cancelled?

OpenStudy (jhannybean):

...

OpenStudy (jhannybean):

It wouldn't get cancelled.

OpenStudy (anonymous):

confused

OpenStudy (jhannybean):

Oh, that \(\pi\), the one inside , yeah that would.

OpenStudy (jhannybean):

\[C= 2\pi \sqrt{\frac{2704\color{red}{\pi}}{\color{red}{\pi}}} \]

OpenStudy (jhannybean):

yeah those do. Sorry.

OpenStudy (anonymous):

so, 2pi*52?

OpenStudy (jhannybean):

and what is \(52 \cdot 2 = \) ?

OpenStudy (anonymous):

104

OpenStudy (jhannybean):

There you go.

OpenStudy (anonymous):

there were 3 pi's i guess

OpenStudy (jhannybean):

Yeah, and you were left with 1.

OpenStudy (anonymous):

so it's104cm?

OpenStudy (jhannybean):

\[C=104\pi\]

OpenStudy (jhannybean):

Don't forget the \(\pi\).

OpenStudy (jhannybean):

\[C= 2\pi \sqrt{\frac{2704\color{red}{\pi}}{\color{red}{\pi}}} = 2\pi \cdot 52 = 104\pi \approx 326.7256\]

OpenStudy (anonymous):

one more question.

OpenStudy (anonymous):

thx for the help

OpenStudy (jhannybean):

And no problem.

OpenStudy (anonymous):

find the circumference of the circle whose area is 16times the area of the cirlce with diameter 1.4cm

OpenStudy (anonymous):

circle*

OpenStudy (jhannybean):

Would you mind closing this question and posting up a new one? :)